1/2

Ирина Саверьевна Алексеева ,  Ирина Саверьевна Алексеева ,
Ode a Samarcanda  Ode to Samarkand
tradukita de Carlo Minnaja  tradukita de Malcolm Jones

Argiope-informo: La aspekto de tiu ĉi versio estas barita. Supozeble la aŭtoro decidis tion.

I sing of you, oh, Samarkand’! Your wondrous features win my praises. And your dear heart - the Registan - seems young to me and never ages.

Its steady beating Ulughbek could sense beneath the constellations. Oh, Samarkand! Immortal pick of mighty Tamerlan’s creations.

Madrasses are your sacred schools, which cultivate the arts and learning. Their beauty overcomes time’s rules by centuries into moments turning.

Here the great Navoi’s verse, so full of love and strong emotion, is cherished by your populace, made happy by its inspiration.

The domes of minarets stand proud - like day-time stars in clear sky shining: so pure the blue, without a cloud and fine the spires in art combining.

Oh, Samarkand! You’re wise through age, with mosques and churches closely neighboured. For those who love you, the world’s stage has no town better nor more favoured.

Oh, Samarkand! Live long and thrive, be glorified by each good action! And while eternity’s your prize, be here our lasting benefaction.

…  …

2/2

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en itala de Carlo Minnaja (∗1940-03-19).

Arg-860-2206 (2014-11-13 12:31:20)

Tiu ĉi poem-italigo aperis en 2013 en la libro “Oда Самарканду”. Ода Самарканду: пер. на 62 языка / И. С. Алексеева ; ред.-сост.: А. Ионесов . – Москва. : Изд-во журн. “Юность”, 2013 . – 91 с. : ил., 2 л. ил. Pri la tradukinto vidu la retejon http: // it. wikipedia. org/ wiki/ Carlo_Minnaja .

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en la Anglan de Malcolm Jones.

Arg-860-2037 (2014-10-16 20:52:12)

Tiu ĉi poem-angligo aperis en 2013 en libro “Oда Самарканду” (= Odo por Samarkando). Sinjoro Malcolm Jones, kiu angligis la poemon, estas brita esperantisto. Lia adreso troviĝas en la UEA-jarlibro de la jaro 2014 sur paĝo 126.

http://www.poezio.net/version?poem-id=860&version-id=2206,2037